Suppose that your car consumes a gallon of gasoline in half an hour. What is its rate of chemical energy consumption, in watts and in horsepower?


1
Expert's answer
2021-11-30T17:44:35-0500

A gallon of gasoline contains 1.3·108 J of energy. In 30 minutes, or 30·60=1800 seconds, this is equivalent to the power of


"P=\frac{E}{t}=\frac{1.3·10^8}{1800}=72.2·10^3\text{ W}."

Since 103 kW is 1.34 hp, this is equivalent to


"P=72.2·1.34=96.7\text{ hp}."